# Ruminations Upon the Modeling of X-ray Foregrounds, Backgrounds and Faint Sources

### v1

@@ Abstract
1 Introduction
If you want to study the X-ray emission from clusters of galaxies, you will end up actually studying the X-ray background. – Apocryphally attributed to S. Snowden
@@ 1 Introduction
Direct application of such empirical background estimates is also problematic from a statistical perspective. The use of a Gaussian likelihood function to fit fundamentally Poisson data biases results, even when binning the data to hundreds of counts or more (Cash, 1979; Wachter et al., 1979; Nousek & Shue, 1989; Leccardi & Molendi, 2007; Humphrey et al., 2009; Mantz et al., 2017; Bonamente, 2020). Use of a Poisson likelihood (equivalently, the Cash fitting statistic; Cash 1979) ameliorates this issue but, unlike a Gaussian likelihood, it does not provide a simple way to directly subtract a background estimated with a given (also Gaussian) uncertainty. A common solution in spectroscopy is to use the modified Cash statistic which approximately marginalizes over the background signal in each channel, as constrained by a separate, empirical and similarly Poisson measurement of the background. [^2] https://heasarc.gsfc.nasa.gov/docs/software/xspec/manual/node336.html This approach produces unbiased measurements in practice (Humphrey et al., 2009; Mantz et al., 2017; Bonamente, 2020), though it still requires data to be binned to $\geq 1$ count to do so. [^3] We are not aware of a refereed citation for this fact, but it can be straightforwardly verified by simulation. While this may not seem like a great imposition, it can be a limitation for imaging-spectroscopic applications, requiring one to compromise either spatial or spectral resolution unnecessarily.
Figure 1: Visualization of a particular Chandra blank-sky spectrum for a $2^{\prime\prime}$ radius circle, compared with the corresponding generative model. Green points (corresponding to individual events) show the unbinned spectrum. Blue crosses show the spectrum grouped to a minimum of 1 count per bin using the “group min 1” command to the grppha tool. Pink crosses show the same spectrum binned according to the same grouping procedure applied to the corresponding science spectrum. Error bars on these points show the 68.3 percent credible intervals on the spectrum in each bin, assuming uniform and independent priors over the non-negative real line. The black curve shows the model produced by our modification of the Suzuki et al. (2021) procedure (see Section 3.3.1).
@@
In short, with the community’s accumulated 25 years of experience with the instrumental backgrounds in Chandra and XMM, and concurrent advances in our understanding of the astrophysical backgrounds, the time seems right to ask whether a complete generative model of these unwanted signals can now be employed, and what affect this might have on measurements of the ICM compared with other approaches. This paper describes methods we have developed for defining and constraining such background models and applying them to particular Chandra or XMM observations. These are in no sense the uniquely optimal answer to X-ray background modeling, but provide a useful benchmark. We demonstrate our approach with analyses of several galaxy clusters, spanning a range of redshift, mass and data quality. The results are compared with those obtained using direct, empirical estimation of the backgrounds from blank-sky observations or source-free regions. More prosaically, this paper serves as a reference for the methodology we are applying in a number of ongoing X-ray cluster studies.
@@
Section 2 reviews the standard data reduction procedures we employ for Chandra and XMM. In Section 3, we describe the models of the various background components, while Section 4 covers the procedure for applying them to a particular target and set of observations. Section 5 presents multiple example analyses with somewhat different goals, to demonstrate the impact of these methods. We conclude in Section 6.
2 Data reduction
@@
This section summarizes the data reduction process for both Chandra and XMM. Specific observations analyzed for this work are introduced on a per-cluster basis in Section 5. The Chandra data are compiled in Chandra Data Collection 491, [^5] https://doi.org/10.25574/cdc.491 (catalog doi:10.25574/cdc.491) while XMM data can be downloaded from the XMM-Newton Science Archive. [^6] https://www.cosmos.esa.int/web/xmm-newton/xsa
2.1 Chandra
@@
Our procedure for reducing the Chandra data follows the recommendations in the Chandra Analysis Guide [^7] https://cxc.harvard.edu/ciao/guides/acis_data.html and was previously described by Mantz et al. (2015). Apart from using newer versions of the Chandra analysis software (ciao 4.16) and calibration files (caldb 4.11.0), these steps remain unchanged. In brief, level-2 event files are regenerated from the level-1 data using the ciao script acis_process_events, including standard grading and the use of VFAINT mode information where available. We then use the ciao tool wavdetect to perform an initial identification of point-like sources, which may be time-variable AGN, and mask them while extracting light curves used to filter periods of enhanced background. This filtering is generally accomplished using the lc_clean tool with default settings, except in rare cases where human intervention is required (typically when an extended background flare biases the mean rate even after sigma clipping). When multiple observations of the same target exist, we also use the preliminary point source detections above to correct their relative astrometry. This is of minimal importance for measurements of the intracluster medium, given Chandra’s overall good pointing accuracy, but ensures that our final detection of point-like sources (Section 3.2) is as accurate as possible.
@@ 2.2 XMM-Newton
3 Modeling of foregrounds, backgrounds and the ICM
[…]
3.1 Galactic X-ray foreground
The soft X-ray foreground originating in the local Universe is typically described as the sum of emission from the local hot bubble and Galactic halo, potentially including contributions from unresolved stellar and extragalactic sources. Modeling based on ROSAT data found that the typical foreground spectrum could be described as the sum of 2 thermal emitters with temperatures $\sim 100$ eV (one absorbed by Galactic hydrogen and one unabsorbed, reflecting a distribution in the distance to the emitters) and a hotter, absorbed thermal component with a typical temperature of $\sim 200$ eV (e.g. Snowden et al. 1998; Kuntz & Snowden 2000). With the higher spectral resolution of XMM and Suzaku, additional contributions from SWCX (e.g. Snowden et al. 2004; Koutroumpa et al. 2009; Ueda et al. 2022) and yet hotter thermal emission at $\sim 700$ eV (e.g. Kuntz & Snowden 2008; Das et al. 2019, 2021; Huang et al. 2023) have been discussed.
@@ 3.1 Galactic X-ray foreground
3.1.1 Chandra
Since Chandra currently has relatively little sensitivity at soft energies, [^9] Based on the data for A1795 that we analyze in Appendix A, contamination on the optical blocking filter (e.g. Plucinsky et al. 2003, 2018) has reduced the sensitivity of ACIS-S and ACIS-I since launch by factors of $\sim 100$ and 10, respectively. we obtain an independent constraint from ROSAT All-Sky Survey (RASS) data from nearby regions of the sky. Specifically, we used the heasoft tool sxrbg to obtain average X-ray background count rates across the six bands of the RASS diffuse background maps (Snowden et al., 1997; Sabol & Snowden, 2019) in an annulus between $0.5^{\circ}$ and $1.0^{\circ}$ from the nominal cluster position (these radii would need to be increased for particularly extended/low-redshift clusters). Following Sabol & Snowden (2019), [^10] https://heasarc.gsfc.nasa.gov/Tools/xraybg we model the Galactic soft foreground as the sum of an unabsorbed thermal plasma component and two absorbed thermal components, where the absorbing equivalent H I column density is that in the direction of the cluster. The unabsorbed and one absorbed component are assumed to have temperatures of 0.1 keV while the temperature of the third component is free, and all 3 are assumed to have solar metallicity. A power-law component with a photon index of 1.45 is also included in the model, to account for the unresolved (at ROSAT resolution) CXB. The free parameters of the model thus consist of the temperature of one unabsorbed thermal emitter, and normalizations for all 3 thermal components and the power-law model. Apart from the power-law component, which is handled differently in later analysis, the best fit to the ROSAT data defines our fiducial model of the soft foreground.
@@ 3.1.1 Chandra
When analyzing Chandra data, we similarly fix the temperatures and relative normalization of the thermal components, letting only their overall normalization vary. (In practice, the 0.1 keV components contribute so little in the energy band used in this analysis that they could be safely ignored.) Typically, we fix all of the foreground temperature parameters at this stage, and marginalize over the normalization using a Gaussian prior defined by the ROSAT uncertainty, above. In rare cases, the local foreground is bright enough that Chandra can provide comparable or better information on the normalization and/or temperature of the warm component. In Section 4.2, we will assess this and revise the foreground priors if warranted.
3.1.2 XMM
@@
We also allow for the possibility of an additional thermal component with temperature $\sim 0.75$ keV. This component is not included by default, but its presence is tested for when spectrally analyzing regions free of cluster emission (Section 4.2). Given such a region, as always exists in this work, the foreground model is fit to the data without external priors.
3.2 Point-like sources and the CXB
@@
The Chandra-detected point source masks are resized for XMM analysis following the strategy of Mantz et al. (2020). The wavdetect count rate for each source was converted to an approximate 2–10 keV flux assuming a power-law spectrum with a photon index of 1.4, and the radius of the corresponding mask chosen such that the maximum unmasked surface brightness according to the PSF model is $5\times 10^{-19}$ erg s-1 cm-2 arcsec-2 in the 0.4–4.0 keV band (with a maximum radius of $60^{\prime\prime}$; Figure 4). This seemingly arbitrary value was originally chosen to produce masks that conservatively cover the visible emission. The residual flux beyond the masks can nevertheless be computed given the Chandra flux estimates and XMM PSF model, and is typically negligible compared with other backgrounds. In rare cases, we adopt a smaller mask for sources projected onto a cluster in order to avoid masking too much emission from the ICM (Section 5.4; Mantz et al. 2020; Flores et al. 2021); here it naturally is more important to model the residual, unmasked emission. In addition to the Chandra-detected sources, we visually identify and mask (to a comparable surface brightness) sources in the XMM data that were too faint to be detected at the time of the Chandra observations or fell outside of Chandra’s field of view. Note that, in this work, the regions identified for analysis of the ICM are entirely covered by Chandra.
@@
In the context of point source detection, the “background” surface brightness density consists of all of the foreground and background components discussed in this section (including all other undetected point sources) and the ICM or other target of observation. An accurate prediction for the unresolved CXB can thus only be made following a preliminary imaging analysis of the observed field, as we perform in Section 4.1. Once such a model of the “background” as a function of position is available, we can compute a corresponding map of the sensitivity limit for point source detections, converting the limiting mean number of counts per pixel to a limiting unabsorbed surface brightness in the 2–10 keV band, assuming a power-law spectrum with photon index 1.7. Following de Vries et al. (2023), we then integrate the AGN luminosity function model of Miyaji et al. (2015) up to this limit as a function of position to form a predicted map of emission from undetected CXB sources. This map can be converted to absorbed surface brightness in an appropriate energy band for subsequent imaging analysis, or integrated in a region of interest to normalize a CXB model used in spectral analysis (Figure 3, bottom). When modeling the unresolved CXB, primarily consisting of sources much fainter than the sensitivity limit, we assume a shallower photon index of 1.4.
3.3 Quiescent particle-induced background
3.3.1 Chandra
@@
3.3.2 XMM
For XMM analysis, filter-closed data perform the same function as the Chandra stowed data sets. Furthermore, sas has long provided “background” data sets tailored to a given observation by resampling filter-closed data taken under comparable conditions, as determined from events recorded in the unexposed “corner” regions of the EPIC detectors, excluding energy ranges associated with certain fluorescence lines. Standard sas/esas processing produces such a tailored background file for each science image and spectrum. However, these background files cannot be directly reinterpreted as a model because they contain Poisson fluctuations from the original filter-closed data that do not represent our expectation for the mean background in a given pixel or channel.
We find that a small amount of Gaussian smoothing, at or below the PSF scale or spectral resolution, is sufficient to regularize the background such that Poisson fluctuations in the original data do not unduly influence the results. In the imaging context, edge effects are typically negligible, provided that unobserved pixels are excluded from the averaging rather than being included with values of zero. For pn spectra, smoothing with a 25 eV (standard deviation) kernel does cause the smoothed spectrum to systematically depart from the original, but only at energies below the minimum of 0.4 keV that we use for analysis.

### v2

@@ Abstract
I Introduction
@@
Direct application of such empirical background estimates is also problematic from a statistical perspective. The use of a Gaussian likelihood function to fit fundamentally Poisson data biases results, even when binning the data to hundreds of counts or more (Cash, 1979; Wachter et al., 1979; Nousek & Shue, 1989; Leccardi & Molendi, 2007; Humphrey et al., 2009; Mantz et al., 2017; Bonamente, 2020). Use of a Poisson likelihood (equivalently, the Cash fitting statistic; Cash 1979) ameliorates this issue but, unlike a Gaussian likelihood, it does not provide a simple way to directly subtract a background estimated with a given (also Gaussian) uncertainty. A common solution in spectroscopy is to use the modified Cash statistic which approximately marginalizes over the background signal in each channel, as constrained by a separate, empirical and similarly Poisson measurement of the background. [^2] https://heasarc.gsfc.nasa.gov/docs/software/xspec/manual/node336.html This approach produces unbiased measurements in practice (Humphrey et al., 2009; Mantz et al., 2017; Bonamente, 2020), though it still requires data to be binned to $\geq 1$ count to do so. [^3] We are not aware of a refereed citation for this fact, but it can be straightforwardly verified by simulation. While this may not seem like a great imposition, it can be a limitation for imaging-spectroscopic applications, requiring one to compromise either spatial or spectral resolution unnecessarily. Binning aside, the notions that we have so little information about the background (i.e., no more than is present in the background region of a particular observation) and that the background in every channel is independent of every other strain credulity (see also discussions by van Dyk et al. 2001; Broos et al. 2010).
Figure 1: Visualization of a particular Chandra blank-sky spectrum for a $2^{\prime\prime}$ radius circle, compared with the corresponding generative model. Green points (corresponding to individual events) show the unbinned spectrum. Blue crosses show the spectrum grouped to a minimum of 1 count per bin using the “group min 1” command to the grppha tool. Pink crosses show the same spectrum binned according to the same grouping procedure applied to the corresponding science spectrum. Error bars on these points show the 68.3 percent credible intervals on the spectrum in each bin, assuming uniform and independent priors over the non-negative real line. The black curve shows the model produced by our modification of the Suzuki et al. (2021) procedure (see Section III.3.1).
@@
In short, with the community’s accumulated 25 years of experience with the instrumental backgrounds in Chandra and XMM, and concurrent advances in our understanding of the astrophysical backgrounds, the time seems right to ask whether a complete generative model of these unwanted signals can now be employed, and what affect this might have on measurements of the ICM compared with other approaches. This paper describes methods we are employing in concurrent work for defining and constraining such background models and applying them to particular Chandra or XMM observations. These are not necessarily novel individually, and similarly motivated modeling of various background components can be found throughout the literature (e.g., Snowden et al. 2008; Eckert et al. 2014; Sanders et al. 2022). Indeed, our purpose here is not primarily to contribute substantively to the physical modeling of any of these backgrounds, but to synthesize the best practices for modeling them that are currently practical, in order to provide a useful benchmark. In particular, few have taken advantage of the QPB models for ACIS produced by Bartalucci et al. (2014) and Suzuki et al. (2021), and the practice of simultaneously fitting complete background and source models, as opposed to estimating and subsequently fixing background models, remains rare. We demonstrate our approach with analyses of several galaxy clusters, spanning a range of redshift, mass and data quality, and compare the results with those obtained using direct, empirical estimation of the backgrounds from blank-sky observations or source-free regions. More prosaically, this paper serves as a reference for the methodology we are applying in a number of ongoing X-ray cluster studies.
@@
Section II reviews the standard data reduction procedures we employ for Chandra and XMM. In Section III, we describe the models of the various background components, while Section IV covers the procedure for applying them to a particular target and set of observations. Section V presents multiple example analyses with somewhat different goals, to demonstrate the impact of these methods. We conclude in Section VI.
II Data reduction
@@
This section summarizes the data reduction process for both Chandra and XMM. Specific observations analyzed for this work are introduced on a per-cluster basis in Section V. The Chandra data are compiled in Chandra Data Collection 491, [^5] https://doi.org/10.25574/cdc.491 (catalog doi:10.25574/cdc.491) while XMM data can be downloaded from the XMM-Newton Science Archive. [^6] https://www.cosmos.esa.int/web/xmm-newton/xsa
II.1 Chandra
@@
Our procedure for reducing the Chandra data follows the recommendations in the Chandra Analysis Guide [^7] https://cxc.harvard.edu/ciao/guides/acis_data.html and was previously described by Mantz et al. (2015). Apart from using newer versions of the Chandra analysis software (ciao 4.16) and calibration files (caldb 4.11.0), these steps remain unchanged. In brief, level-2 event files are regenerated from the level-1 data using the ciao script acis_process_events, including standard grading and the use of VFAINT mode information where available. We then use the ciao tool wavdetect to perform an initial identification of point-like sources, which may be time-variable AGN, and mask them while extracting light curves used to filter periods of enhanced background. This filtering is generally accomplished using the lc_clean tool with default settings, except in rare cases where human intervention is required (typically when an extended background flare biases the mean rate even after sigma clipping). The tool is run on data from a CCD that lacks strong cluster emission, since the mean absolute rate is assumed to reflect the background, generally chip I0 or S1. When multiple observations of the same target exist, we also use the preliminary point source detections above to correct their relative astrometry. This is of minimal importance for measurements of the intracluster medium, given Chandra’s overall good pointing accuracy, but ensures that our final detection of point-like sources (Section III.2) is as accurate as possible.
@@ II.2 XMM-Newton
III Modeling of foregrounds, backgrounds and the ICM
This section covers the models we use to describe each background component, as well as the ICM. A few commonalities can be mentioned at the outset. We use xspec to evaluate each of the spectral models, as well as to compare them to the data. Within xspec, thermal emission is evaluated using the apec model, with metal abundances scaled relative to the Solar table of Asplund et al. (2009). Photoelectric absorption is evaluated using the phabs model, with the equivalent absorbing hydrogen column density in a given direction provided by the HI4PI survey (HI4PI Collaboration, 2016). [^8] Though the issue does not arise in this work, the neutral hydrogen-based estimates of the column density are known to be inaccurate for particularly dense lines of sight ($N_{H}\gtrsim 10^{21}$ cm-2; e.g. Willingale et al. 2013); we recommend allowing the column density parameter to freely fit in such cases. Additional spectral models with more specific purposes will be mentioned below. For both imaging and spectroscopy, we account for the XMM PSF using the symmetric Gaussian+King model developed by Read et al. (2011). Whenever possible, we conservatively exclude the edges of the ACIS CCDs, whose exposure is affected by the Chandra dithering pattern, with masks extending $25^{\prime\prime}$ inward from the nominal chip edges in sky coordinates. Finally, as described in Section III.2, we limit Chandra analysis to chips I0–3 for ACIS-I observations and S1–3 for ACIS-S observations.
III.1 Galactic X-ray foreground
The soft X-ray foreground originating in the local Universe is typically described as the sum of emission from the local hot bubble and Galactic halo, potentially including contributions from unresolved stellar and extragalactic sources. Modeling based on ROSAT data found that the typical foreground spectrum could be described as the sum of 2 thermal emitters with temperatures $\sim 100$ eV (one absorbed by Galactic hydrogen and one unabsorbed, reflecting a distribution in the distance to the emitters) and a hotter, absorbed thermal component with a typical temperature of $\sim 200$ eV (e.g. Snowden et al. 1998; Kuntz & Snowden 2000). With the higher spectral resolution of XMM and Suzaku, additional contributions from SWCX (e.g. Snowden et al. 2004; Koutroumpa et al. 2009; Ueda et al. 2022) and yet hotter thermal emission at $\sim 700$ eV (e.g. Kuntz & Snowden 2008; Das et al. 2019, 2021; Huang et al. 2023) have been discussed. The applicability of this set of model components has been broadly confirmed by eROSITA All-Sky Survey measurements (Ponti et al., 2023; Yeung et al., 2024).
@@ III.1 Galactic X-ray foreground
III.1.1 Chandra
Since Chandra currently has relatively little sensitivity at soft energies, [^9] Based on the data for A1795 that we analyze in Appendix A, contamination on the optical blocking filter (e.g. Plucinsky et al. 2003, 2018) has reduced the sensitivity of ACIS-S and ACIS-I since launch by factors of $\sim 100$ and 10, respectively. we obtain an independent constraint from ROSAT All-Sky Survey (RASS) data from nearby regions of the sky. Specifically, we used the heasoft tool sxrbg to obtain average X-ray background count rates across the six bands of the RASS diffuse background maps (Snowden et al., 1997; Sabol & Snowden, 2019) in an annulus between $0.5^{\circ}$ and $1.0^{\circ}$ from the nominal cluster position (these radii would need to be increased for particularly extended/low-redshift clusters; Figure 2). Following Sabol & Snowden (2019), [^10] https://heasarc.gsfc.nasa.gov/Tools/xraybg we model the Galactic soft foreground as the sum of an unabsorbed thermal plasma component and two absorbed thermal components, where the absorbing equivalent H I column density is that in the direction of the cluster. The unabsorbed and one absorbed component are assumed to have temperatures of 0.1 keV while the temperature of the third component is free, and all 3 are assumed to have solar metallicity. A power-law component with a photon index of 1.45 is also included in the model, to account for the unresolved (at ROSAT resolution) CXB. The free parameters of the model thus consist of the temperature of one unabsorbed thermal emitter, and normalizations for all 3 thermal components and the power-law model. Apart from the power-law component, which is handled differently in later analysis, the best fit to the ROSAT data defines our fiducial model of the soft foreground.
@@ III.1.1 Chandra
When analyzing Chandra data, we similarly fix the temperatures and relative normalization of the thermal components, letting only their overall normalization vary. (In practice, the 0.1 keV components contribute so little in the energy band used in this analysis that they could be safely ignored.) Typically, we fix all of the foreground temperature parameters at this stage, and marginalize over the normalization using a Gaussian prior defined by the ROSAT uncertainty, above. In rare cases, the local foreground is bright enough that Chandra can provide comparable or better information on the normalization and/or temperature of the warm component. In Section IV.2, we will assess this and revise the foreground priors if warranted.
III.1.2 XMM
@@
We also allow for the possibility of an additional thermal component with temperature $\sim 0.75$ keV. This component is not included by default, but its presence is tested for when spectrally analyzing regions free of cluster emission (Section IV.2). Given such a region, as always exists in this work, the foreground model is fit to the data without external priors.
III.2 Point-like sources and the CXB
@@
The Chandra-detected point source masks are resized for XMM analysis following the strategy of Mantz et al. (2020). The wavdetect count rate for each source was converted to an approximate 2–10 keV flux assuming a power-law spectrum with a photon index of 1.4, and the radius of the corresponding mask chosen such that the maximum unmasked surface brightness according to the PSF model is $5\times 10^{-19}$ erg s-1 cm-2 arcsec-2 in the 0.4–4.0 keV band (with a maximum radius of $60^{\prime\prime}$; Figure 4). This seemingly arbitrary value was originally chosen to produce masks that conservatively cover the visible emission. The residual flux beyond the masks can nevertheless be computed given the Chandra flux estimates and XMM PSF model, and is typically negligible compared with other backgrounds. In rare cases, we adopt a smaller mask for sources projected onto a cluster in order to avoid masking too much emission from the ICM (Section V.4; Mantz et al. 2020; Flores et al. 2021); here it naturally is more important to model the residual, unmasked emission. In addition to the Chandra-detected sources, we visually identify and mask (to a comparable surface brightness) sources in the XMM data that were too faint to be detected at the time of the Chandra observations or fell outside of Chandra’s field of view. Note that, in this work, the regions identified for analysis of the ICM are entirely covered by Chandra.
@@
In the context of point source detection, the “background” surface brightness density consists of all of the foreground and background components discussed in this section (including all other undetected point sources) and the ICM or other target of observation. An accurate prediction for the unresolved CXB can thus only be made following a preliminary imaging analysis of the observed field, as we perform in Section IV.1. Once such a model of the “background” as a function of position is available, we can compute a corresponding map of the sensitivity limit for point source detections, converting the limiting mean number of counts per pixel to a limiting unabsorbed surface brightness in the 2–10 keV band, assuming a power-law spectrum with photon index 1.7. Following de Vries et al. (2023), we then integrate the AGN luminosity function model of Miyaji et al. (2015) up to this limit as a function of position to form a predicted map of emission from undetected CXB sources. This map can be converted to absorbed surface brightness in an appropriate energy band for subsequent imaging analysis, or integrated in a region of interest to normalize a CXB model used in spectral analysis (Figure 3, bottom). When modeling the unresolved CXB, primarily consisting of sources much fainter than the sensitivity limit, we assume a shallower photon index of 1.4.
III.3 Quiescent particle-induced background
III.3.1 Chandra
@@
III.3.2 XMM
For XMM analysis, filter-closed data perform the same function as the Chandra stowed data sets. Furthermore, sas has long provided “background” data sets tailored to a given observation by resampling filter-closed data taken under comparable conditions, as determined from events recorded in the unexposed “corner” regions of the EPIC detectors, excluding energy ranges associated with certain fluorescence lines. Standard sas/esas processing produces such a tailored background file for each science image and spectrum. These background files do not constitute a generative model in the sense that we advocate for in Section I, in contrast to the ACIS QPB model of Suzuki et al. (2021), which synthesizes the available background-only data to provide a physically motivated prediction for the background in a given observation. Nevertheless, with the few compromises described below, we can treat them as such in practice.
The SAS background files retain Poisson fluctuations from the original filter-closed data that do not represent our expectation for the mean background in a given pixel or channel. We find that a small amount of Gaussian smoothing, at or below the PSF scale or spectral resolution, is sufficient to regularize the background such that Poisson fluctuations in the original data do not unduly influence the results. In the imaging context, edge effects are typically negligible, provided that unobserved pixels are excluded from the averaging rather than being included with values of zero. For pn spectra, smoothing with a 25 eV (standard deviation) kernel does cause the smoothed spectrum to systematically depart from the original, but only at energies below the minimum of 0.4 keV that we use for analysis.